OpenStudy (anonymous):

Find four - thirds of 9/2

OpenStudy (anonymous):

\[\frac{ 9 }{ 2 } \div\frac{ 4 }{ 3 } =\frac{ 9 }{ 2 }\times \frac{ 3 }{ 4 }\]